# Einstein Fields: A Neural Perspective to Computational General Relativity

### v1

@@ C.0.0.1 Finite-difference method bottlenecks in NR
  • Higher-order finite-difference stencils require a collection of padded grid points exterior to the cube as boundary handling. These are often called ghost cells (zones). For e.g., if $n_{g}$ ghost cells are required for the $n$-th order forward difference stencil, the grid size increases of a 3D spatial voxel grid (Ruchlin et al., 2018) $\mathtt{N}_{x}\times\mathtt{N}_{y}\times\mathtt{N}_{z}\rightarrow(\mathtt{N}_{x}+n_{g})\times(\mathtt{N}_{y}+n_{g})\times(\mathtt{N}_{z}+n_{g})$.
@@
C.0.0.2 Comparing AD vs FD based methods.
@@
Appendix D Experimental details
This appendix provides detailed experimental specifics, including: (i) gradient alignment aspects relevant to SOAP optimizers; (ii) error plots for the metric and higher-rank differential geometric quantities, with component-wise tomography where applicable; (iii) training across varied coordinate systems to illustrate the coordinate-choice flexibility of NeFs; (iv) the hyperparameter configurations employed; and (v) the hardware and software environments used for these experiments.
D.1 Gradient alignment
@@
Competing tasks is a well-known problem in multi-objective learning (Yu et al. (2020), Liu et al. (2021), Shi et al. (2023)), the gradients of the loss functions pull the weights in different directions. In Scientific Machine Learning (SciML), a lot of work emerged in analyzing and mitigating gradient conflicts in the context of PINNs (Wang et al. (2025), Lui et al. (2025), Hwang & Lim (2025)). Although Sobolev training differs from PINNs, particularly from a supervision perspective, it exhibits the same problem where some loss terms dominate others. In PINNs, the training is highly dependent on first satisfying the initial/boundary conditions, which provide uniqueness to the solution. The different levels of complexity between these and the residual loss create different optimization priorities, but both losses are equally important. Similarly, Sobolev training faces analogous challenges with competing loss components. The Jacobian data serve to constrain the model’s derivatives, while the target function outputs determine the integration "constant", both components being equally valuable. However, the sources of complexity differ between these approaches. In PINNs, the primary challenge stems from determining a solution through unsupervised learning on PDE losses, whereas in our Sobolev training specifically, the complexity arises from managing optimization stability in high-dimensional spaces: a $\mathtt{16}$-dimensional output space, $\mathtt{64}$-dimensional Jacobian, and $\mathtt{256}$-dimensional Hessian. Moreover, this complexity is accompanied by the challenge of handling gradient imbalances. Depending on the point in spacetime, the metric or its derivatives dominate in the loss. Generally speaking, an analogy is to think $g_{\mu\nu}\propto\frac{1}{r}$,$\partial_{\rho}g_{\mu\nu}\propto\frac{1}{r^{2}}$ and $\partial_{\sigma\rho}g_{\mu\nu}\propto\frac{1}{r^{3}}$, making it clear how gradient magnitudes differ depending on the radius.
@@
Mitigating gradient conflicts does not necessarily result in better accuracy, but it explains a possible reason why the loss does not improve further, the optimization being stuck in the local minimum of one of the objectives. The intra-step gradient alignment scores presented in Wang et al. (2025) demonstrate SOAP as a far superior alternative compared to other well-established optimizers, or at least for the experiments considered in that study. To provide a direct comparison using the same methodology as in Wang et al. (2025), we evaluated both ADAM and SOAP on the Cartesian KS representation of the Kerr metric, chosen as the most complex metric investigated in this work. The Sobolev training contains only two objectives: metric and Jacobian supervision. For our experimental setup, we employed an MLP architecture with $\mathtt{5}$ hidden layers, $\mathtt{190}$ hidden units per layer, and SILU activation function to compare gradient conflicts between the two optimizers. The training utilized a cosine decay learning rate schedule, starting from an initial learning rate of $\mathtt{1E{-2}}$ and decaying to a final rate of $\mathtt{1E{-8}}$ over $\mathtt{200}$ epochs. For weighting the losses, gradient normalization was used with and without an exponential moving average. As shown in Figure 15, even though Adam is providing twice as much gradient alignment score in almost all epochs, SOAP’s second-order and preconditioning capabilities allow for a 100x training loss improvement.
@@
D.2 Kerr metric error visualization
@@
Figure 17: Kerr metric absolute error between ground truth (analytic) metric and the EinFields parametrized metric. The metrics are depicted in the Cartesian Kerr-Schild (KS) representation as presented in Equation (96). The 2D slice of all the metric components captured in the x-y plane at fixed $z=1.4$ for a spin parameter value $a=0.7$.
D.3 Tomography: Metric, metric Jacobian and metric Hessian components
The effect of introducing losses pertaining to metric Jacobian and Hessian supervision, apart from the metric loss that EinFields predominantly uses, can be quantified and visualized with the following plots below. Here, for the sake of visualization, we do a tomography (2D cuts) of different metric components along a particular axis for the Kerr metric in Cartesian KS coordinates (Equation (96)).
D.4 Training on varied coordinate systems
Understanding this behavior is essential for developing robust INR-based frameworks for representing geometric quantities in numerical relativity, while respecting the underlying diffeomorphism invariance of general relativity.
For the Schwarzschild case, we initiate the training by sampling query spacetime coordinates in spherical representation $(t,r,\theta,\phi)$. These sampled collocation points are then transformed into their corresponding collocation points in Cartesian coordinates $(t,x,y,z)$ and ingoing Eddington-Finkelstein coordinates $(v,r,\theta,\tilde{\phi})$ (see Section B.1.1 for explicit transformation details). Subsequently, EinFields outputs the metric tensors corresponding to these coordinate systems, yielding Equations ((85), (89), (92)).
For the Kerr metric, which is characterized by its oblate spheroidal geometry, we sample query collocation points in the Boyer-Lindquist coordinates $(t,r,\vartheta,\phi)$, followed by the collocation points transformed into Cartesian $(t,x,y,z)$ and ingoing Eddington-Finkelstein coordinates $(t,r,\theta,\tilde{\phi})$. EinFields then outputs the Kerr metric tensors in these respective coordinate systems, resulting in Equations ((95), (96), (98)).

### v2

@@ C.0.0.1 Finite-difference method bottlenecks in NR
  • Higher-order finite-difference stencils require a collection of padded grid points exterior to the cube as boundary handling. These are often called ghost cells (zones). For e.g., if $n_{g}$ ghost cells are required for the $n$-th order forward difference stencil, the grid size increases of a 3D spatial voxel grid (Ruchlin et al., 2018), $\mathtt{N}_{x}\times\mathtt{N}_{y}\times\mathtt{N}_{z}\rightarrow(\mathtt{N}_{x}+n_{g})\times(\mathtt{N}_{y}+n_{g})\times(\mathtt{N}_{z}+n_{g})$.
@@
Appendix D Succinct Introduction to General Relativity, Equations of Motion and Exact Solutions
D.0.0.1 Derivative operators.
D.0.0.2 Differential geometric objects.
[…]
Figure 10: Differential geometry workflow in general relativity (only lhs of the EFEs – Eq. (2)): We describe each quantity starting left: The metric tensor $g_{\alpha\beta}$ defines the spacetime geometry. Its partial derivatives $\partial$ yield the Christoffel symbols $\Gamma^{\gamma}_{\alpha\beta}$, which describe the notion of parallel transport and defines a covariant derivative operation $\nabla_{\alpha}=\partial_{\alpha}+\Gamma_{\alpha}$. The connection also defines the Lie derivative $\mathcal{L}_{v}$ along vector fields $v$. The connection’s derivatives $\nabla$ give the Riemann curvature tensor $R^{\delta}_{\alpha\beta\gamma}$, which encodes tidal forces. The contraction operator $\mathscr{C}=\text{Tr}_{g}$ contracts with the metric, producing the trace part, i.e., the Ricci tensor $R_{\alpha\beta}$. Its subsequent contraction yields the Ricci scalar $R$. The Riemann tensor further splits into the Weyl tensor $C_{\alpha\beta\gamma\delta}$ (trace-free curvature part) and satisfies the algebraic and differential Bianchi identities $R_{[\alpha\beta\gamma]\sigma}=0$ and $\nabla_{[\sigma}R_{\alpha\beta]\gamma\delta}=0$. Together, these geometric objects form the backbone of general relativity, ultimately entering the Einstein field equations through the Einstein tensor $G_{\alpha\beta}$.
D.0.0.3 Conservation laws.
It follows from the contracted Bianchi identities, i.e., cyclic sum of Riemann curvature tensor covariant derivatives (II Bianchi identity – see Eq. (59b)) vanishes identically:
This is a geometric identity that holds for any (torsion-free) connection compatible with the metric. The identity above consequently leads to the covariant derivative of the stress-energy tensor vanishing, that is, $\nabla_{\beta}T^{\alpha\beta}:=0$ (see Eq. (66)), which corresponds to the energy-momentum conservation in general relativity. If required, conservation laws typically feature as soft constraints in PDEs, and are relevant especially when matter distribution/fields are considered.
D.0.0.4 Equations of motion.
$\frac{d^{2}x^{\mu}}{d\tau^{2}}+\Gamma^{\mu}_{\rho\sigma}\frac{dx^{\rho}}{d\tau}\frac{dx^{\sigma}}{d\tau}=0\ .$
$\frac{D^{2}S^{\mu}}{D\tau^{2}}=R^{\mu}_{\alpha\beta\gamma}X^{\alpha}X^{\beta}S^{\gamma}\ ,$
where, $X^{\alpha}$ is a vector field and $\frac{D}{D\tau}=X^{\alpha}\nabla_{\alpha}$ denotes the directional covariant derivative (see Definition 2). Thus, it encodes information about the tidal forces of gravitation.
D.0.1 Analytical (exact) solutions
Exact solutions of the EFEs are metric tensor fields $g_{\alpha\beta}$ that satisfy Eq. (2). Many exact solutions are known, which can be classified into exterior (vacuum) solutions and interior solutions (Misner et al., 2017). While there exist several geometries that satisfy the EFEs, we shall focus on three prominent vacuum solutions: the Schwarzschild metric, the Kerr metric, and gravitational waves. These geometries not only have a high theoretical and historical relevance, but are also of great interest in computational black hole astrophysics and gravitational wave and multi-messenger astronomy. Appendix B discusses these solutions in more detail, including other prominent coordinates, as well as real and fake (coordinate) singularities. From here on, we work in naturalized units by setting $c=G=1$.
D.0.1.1 Schwarzschild metric
$ds^{2}=-\bigg(1-\frac{2M}{r}\bigg)dt^{2}+\bigg(1-\frac{2M}{r}\bigg)^{-1}dr^{2}+r^{2}\big(d\theta^{2}+\text{sin}^{2}\theta d\phi^{2}\big)\ .$
D.0.1.2 Kerr metric
The metric can be described in the corresponding oblate spheroidal coordinates also known as the Boyer-Lidquist (BL) coordinates $(t,r,\vartheta,\phi)$ – see Eq. (78(Boyer & Lindquist, 1967):
D.0.1.3 Linearized gravity
$h_{\alpha\beta}^{\text{TT}}=\begin{pmatrix}0&0&0&0\\ 0&h_{+}&h_{\times}&0\\ 0&h_{\times}&h_{+}&0\\ 0&0&0&0\end{pmatrix}\cos\big(\omega(t-z)\big)\ .$
Here, $h_{+}$ and $h_{\times}$ are the amplitudes of the “++” (plus) polarization and $\times$” (cross) polarization
Appendix E Related work continued
E.0.0.1 Compression techniques in scientific computing.
classical compression strategies have been a versatile tool in reducing data sizes of large-scale numerical simulation data, which constitute storage memory-intensive domain. Simulation runs can range between petabytes to exabytes of data, thus requiring compression strategies to be integrated within the simulation pipeline, for efficient data volume reduction. These range from lossless and lossy compression techiques (Lindstrom & Isenburg, 2006; Di & Cappello, 2016; ISO Central Secretary, 2024) for multidimensional weather modeling tasks (Huang et al., 2016; 2025), discrete wavelet tranform (DWT)–based compression (Rho et al., 2023) or multidimensional data via tensor decomposition (Ballester-Ripoll et al., 2020).
Recently increasing works have leveraged implicit neural representations for a lossy neural compression (Dupont et al., 2021) by embedding high-dimensional (explicit grids), time-dependent physical simulations into compact, differentiable network weights. Such representations achieve compression factors of several orders of magnitude while retaining physical accuracy and offering efficient gradient access for downstream analysis or control. Applications include multidimensional weather and climate modeling (Huang & Hoefler, 2023) or even hybrid compression techniques using turbulent plasma–based simulations (Galletti et al., 2025).
E.0.0.2 ML applied to gravitational physics.
Gravitational wave modeling and numerical relativity problems have been tackled by state-of-the-art deep learning methods. These include DINGO – a rapid gravitational wave parameter estimation toolkit using NNs as surrogates for Bayesian posterior distributions (Dax et al., 2021). They show orders of magnitude reduction in inference time, bringing it down from $\mathcal{O}(\text{day})$ to $20s$. Similar lines of work include DINGO-BNS that performs real-time inference for binary neutron star (BNS) mergers and applicable for multi-messenger astronomy (Dax et al., 2025). On the other hand, physics informed neural networks (PINNs) (Raissi et al., 2019), have found applications in general relativistic phenomena and NR, such as solving the Tekoulsky equation inorder to compute the first quasinormal modes of the blackholes, for e.g., Kerr geometry (Luna et al., 2023; Cornell et al., 2022). Some other works explore physics informed neural operators (PINO) (Rosofsky & Huerta, 2023) for magnetohydrodynamics, or solving vaccum Einstein equations (Hirst et al., 2025)
Physics-informed neural networks, neural operators and neural fields. Physics-Informed Neural Networks (PINNs) augment neural network training with physical constraints derived from governing differential equations (Karniadakis et al., 2021). This is typically achieved by adding loss terms that penalize violations of the residuals of the underlying PDEs or ODEs (Raissi et al., 2019), together with constraints from boundary conditions and conservation laws. PINNs are primarily data-free approaches and have been especially effective for solving forward and inverse problems. While PINNs share with neural fields the use of coordinate-based neural networks, their purpose is fundamentally different: PINNs solve a physical equation by enforcing its residual during optimization, whereas neural fields represent a given physical field directly from data without solving the governing equations. In this sense PINNs can be considered as a special case of neural fields (Xie et al., 2021), that satisfy the govering physical equations at each step.
Neural operators, in contrast, aim to learn mappings between infinite-dimensional function (Banach) spaces (Kovachki et al., 2023), providing a data-driven approach to approximate solution operators for entire families of PDEs. They prioritize generalization across different input functions, geometries, or forcing conditions, typically requiring large training sets covering many PDE instances.
Implicit neural representations (Müller et al., 2022) occupy a distinct position relative to both paradigms. Rather than enforcing a PDE residual (as in PINNs) or learning an operator over families of solutions (as in neural operators), neural fields provide a compact, continuous, and fully differentiable representation of a single high-dimensional physical field. This makes them especially well suited for encoding scientific data with high fidelity, enabling continuous spatial (and temporal) query access, implicit compression, and differentiable downstream analysis. In this sense, neural fields are not a competing method for PDE solution or operator learning, but a complementary representation framework for capturing and reconstructing complex physical domains.nstructing the dynamics with high fidelity, setting them apart from PINNs and neural operators in general.
Appendix F Experimental details
This appendix provides detailed experimental specifics, including: (i) AD as a superior differentiation framework for tensor fields compared to higher-order finite-differencing methods; (ii) Effectivity of higher derivative losses for retrieving high-precision dynamics and higher-order curvature tensors & invariants; (iii) setup – data preparation (iv) gradient alignment aspects relevant to SOAP optimizers; (v) Component-wise tomography of compression error on the implicit metric and its derived higher-rank differential geometric quantities; (vi) training across varied coordinate systems to illustrate the coordinate-choice flexibility of NeFs; (vii) hyperparameter configurations employed; and (viii) the hardware and software environments used for these experiments.
F.1 Evaluation criteria
These are applied to the metric tenors and their derived quantities, illustrated in Figures 10 and 2. Recall that the tensor components are coordinate-dependent (and even more so, the metric Jacobian, metric Hessian, and Christoffel symbols are not even tensors), and, hence, these errors lack an immediate physical meaning. This is improved with the consideration of scalar quantities such as the Ricci scalar, Kretschmann invariants, and Weyl scalars, which by definition are coordinate-independent quantities.
F.2 Data Generation
Our use cases are exact analytic solutions to the EFEs, i.e., the set of metrics $g_{\alpha\beta}$ that satisfy the Eq. (2). These solutions describe the exterior (vacuum) solutions around massive gravitating objects. For our main set of experiments, we fit a NeF against the analytic solutions introduced in Section D.0.1, each having different features and spatio-temporal symmetries:
  • Schwarzschild metric in spherical coordinates – Eq. (98),
  • Kerr metric in Boyer-Lindquist and Kerr-Schild coordinates – Eqs. (101, 81),
  • gravitational waves metric (TT gauge) in Cartesian coordinates – Eq. (110).
For each, we compute the distortion after subtracting the flat background metric in that particular coordinate chart. Detailed information on data specifications is provided in Table 7.
Additionally, we train each geometry in different coordinate systems to investigate how the choice of coordinates impacts NeFs (recall: the physical laws do not depend on the coordinate system).
Metric Coordinates Domain Resolution Parameters
Schwarzschild Spherical
$(t,r,\theta,\phi)$
$t=0$
$r\in$ [2.5,150]
$\theta\in$ (0, $\pi$)
$\phi\in[0,2\pi)$
1
128
128
128
M = 1
Kerr Boyer-Lindquist
$(t,r,\vartheta,\phi)$
$t=0$
$r\in$ [3,14]
$\vartheta\in$ (0, $\pi$)
$\phi\in$ [0, $2\pi$)
1
128
128
128
M = 1
$a\in$ [0.628,0.95]
Kerr-Schild
$(t,x,y,z)$
$t=0$
$x\in$ [-3, 3]
$y\in$ [-3, 3]
$z\in$ [0.1, 3]
1
128
128
128
M = 1
$a$ = 0.7
Linearized gravity Cartesian
$(t,x,y,z)$
$t\in$ [0, 10]
$x\in$ [0, 10]
$y\in$ [0, 10]
$z\in$ [0, 10]
140
10
10
140
$\omega$ = 1
$\epsilon=10^{-6}$
F.3 Comparing AD vs FD based methods.
@@ F.2 Data Generation
F.4 Higher tensor derivative losses – Sobolev training
where $D^{(j)}$ denotes the $j^{\text{th}}$ derivative operator, which in our case could be the partial derivatives $\partial^{j}$ or covariant derivatives $\nabla^{j}$, and $\lambda_{j}$ are weighting coefficients. This loss promotes alignment not only in function space but also in the Sobolev space $W^{N,2}(\mathcal{X})$, which encodes both value and derivative information. Sobolev training enhances generalization, stability, and accuracy of NeF derivatives (Chetan et al., 2024).
3:for epoch =1=1 to $N_{\text{epochs}}$ do
5:  Move $(x_{\text{batch}},g_{\text{batch}})$ to device $\mathcal{D}$
8:  if $N\geq 1$ then {Jacobian supervision}
9:   Compute $\partial_{x}^{(1)}\hat{g}_{\text{batch}}$ through AD
11:  end if
12:  if $N\geq 2$ then {Hessian supervision}
13:   Compute $\partial_{x}^{(2)}\hat{g}_{\text{batch}}$ through AD
315:  end if
519:end for
22:end for
23:return optimized parameters $\theta$
Algorithm 1 EinField training scheme
[…]
F.5 Gradient alignment
@@
Competing tasks is a well-known problem in multi-objective learning (Yu et al. (2020), Liu et al. (2021), Shi et al. (2023)), the gradients of the loss functions pull the weights in different directions. In Scientific Machine Learning (SciML), a lot of work emerged in analyzing and mitigating gradient conflicts in the context of PINNs (Wang et al. (2025a), Lui et al. (2025), Hwang & Lim (2025)). Although Sobolev training differs from PINNs, particularly from a supervision perspective, it exhibits the same problem where some loss terms dominate others. In PINNs, the training is highly dependent on first satisfying the initial/boundary conditions, which provide uniqueness to the solution. The different levels of complexity between these and the residual loss create different optimization priorities, but both losses are equally important. Similarly, Sobolev training faces analogous challenges with competing loss components. The Jacobian data serve to constrain the model’s derivatives, while the target function outputs determine the integration "constant", both components being equally valuable. However, the sources of complexity differ between these approaches. In PINNs, the primary challenge stems from determining a solution through unsupervised learning on PDE losses, whereas in our Sobolev training specifically, the complexity arises from managing optimization stability in high-dimensional spaces: a $\mathtt{16}$-dimensional output space, $\mathtt{64}$-dimensional Jacobian, and $\mathtt{256}$-dimensional Hessian. Moreover, this complexity is accompanied by the challenge of handling gradient imbalances. Depending on the point in spacetime, the metric or its derivatives dominate in the loss. Generally speaking, an analogy is to think $g_{\mu\nu}\propto\frac{1}{r}$$\partial_{\rho}g_{\mu\nu}\propto\frac{1}{r^{2}}$ and $\partial_{\sigma\rho}g_{\mu\nu}\propto\frac{1}{r^{3}}$, making it clear how gradient magnitudes differ depending on the radius.
@@
Mitigating gradient conflicts does not necessarily result in better accuracy, but it explains a possible reason why the loss does not improve further, the optimization being stuck in the local minimum of one of the objectives. The intra-step gradient alignment scores presented in Wang et al. (2025a) demonstrate SOAP as a far superior alternative compared to other well-established optimizers, or at least for the experiments considered in that study. To provide a direct comparison using the same methodology as in Wang et al. (2025a), we evaluated both ADAM and SOAP on the Cartesian KS representation of the Kerr metric, chosen as the most complex metric investigated in this work. The Sobolev training contains only two objectives: metric and Jacobian supervision. For our experimental setup, we employed an MLP architecture with $\mathtt{5}$ hidden layers, $\mathtt{190}$ hidden units per layer, and SILU activation function to compare gradient conflicts between the two optimizers. The training utilized a cosine decay learning rate schedule, starting from an initial learning rate of $\mathtt{1E{-2}}$ and decaying to a final rate of $\mathtt{1e{-8}}$ over $\mathtt{200}$ epochs. For weighting the losses, gradient normalization was used with and without an exponential moving average. As shown in Figure 15, even though Adam is providing twice as much gradient alignment score in almost all epochs, SOAP’s second-order and preconditioning capabilities allow for a 100x training loss improvement.
@@
F.6 Reconstructing general relativistic dynamics and curvature scalars
We use synthetic data generated from analytical solutions to validate and characterize EinFields. We primarily focus on two interesting aspects of general relativity: (i) general relativistic dynamics, particularly geodesic motion around massive gravitating objects (see Section D.0.0.4) and (ii) global curvature structures encoded in tensorial invariants.
F.6.1 Floating point precision caveat.
F.6.1.1 EinField-based geodesics.
To compute geodesic motion, we numerically integrate the trajectories using a fifth-order explicit singly diagonally implicit Runge-Kutta (ESDIRK) solver. Specifically, we evolve Eq. (56) with respect to the affine parameter $\tau$ (proper time) [^9] Not to be confused with the coordinate time $t$., generating ground-truth geodesics from the analytic Christoffel symbols for the Schwarzschild and Kerr spacetimes. These are then compared against the rollouts obtained using the EinField-reconstructed Christoffel symbols.
To accurately retrace geodesic orbits, it is essential to incorporate Jacobian supervision within the Sobolev training framework. In contrast, additional Hessian supervision results in only marginal improvements for geodesic simulations and is not required in practice. Following Section F.6.1, all geodesic solvers are executed in double precision to ensure numerical stability and high-fidelity trajectory reconstruction.
F.6.2 Schwarzschild metric geodesic setup
$(t,r,\theta,\phi)(t=0)=\big(0,a_{0}r_{s},\pi/2,0\big)$ (106)
and the initial four-velocity
Where $v_{0}=b_{0}\sqrt{1/(r_{0}-r_{s})}$ and $a_{0},b_{0}\in\mathbb{R}$ can be chosen freely to select the desired orbit in the $\theta=\pi/2$ plane. The geodesics in Figure 4 demonstrate a good qualitative agreement over several orbits. The error is quantified and discussed further in Section F.7.2.
F.6.3 Schwarzschild blackhole render
Being able to compute geodesics is sufficient to perform rendering. We use the Schwarzschild EinFields metric to render a black-hole on a celestial background. This requires propagating geodesics from the camera observer via the spacetime terminating at the distant background. The resulting ray-traced image as shown in the main text provides visual evidence for the global consistency and quality of the metric and the derived Christoffel symbols.
F.6.4 Kerr metric geodesic setup
While, the mathematical description of geodesics around Kerr metric is beyond scope for this work (see Teo (2003) for detailed exposition), we consider solving three different cases numerically. These include Zackiger orbits (retrograde geodesics – stable geodesics with larger radii), prograde orbits (stable geodesics with smaller radii), and arbitrary eccentric orbits, which depend on the initial conditions, including choice of energy $E$ and angular momentum $L_{z}$ of the test particle.
F.6.5 Kerr Kretschmann invariant
The Kretschmann invariant (scalar), $\mathscr{K}=R^{\alpha\beta\gamma\delta}(x^{\mu})R_{\alpha\beta\gamma\delta}(x^{\mu}),$ is a key curvature invariant distinguishing true (curvature) singularities from coordinate (apparent) singularities. The nontrivial part of the Kretschmann invariant for the Kerr metric reads:
$r=0\ ,\text{and},\ \vartheta=\dfrac{\pi}{2}\ .$
Thus, the rotation $a$ induces a ring singularity at radius $a$ on the equatorial plane $\theta=\pi/2$, where the curvature diverges. Accurately capturing this geometric structure requires isolating true singularities from coordinate artifacts, which can otherwise lead to incorrect classification of singularities.
We perform training in Cartesian KS coordinates (see Eq. (81)) to eliminate coordinate singularities that would otherwise impede convergence. We first train EinFields (+Jac + Hess) on Cartesian KS coordinates, subsequently constructing the Riemann tensor – see Section A.3.5 via successive automatic differentiation steps and raising indices using the parametrized metric $\hat{g}$ (see Eq. (108)). The NeF reconstructed $\hat{\mathscr{K}}$ Figure 13(b) captures the ring singularity structure and agrees well with the analytical solution, as shown in Figures 13(a). However, the reconstruction remains sensitive to floating-point errors and requires high NeF accuracy for stability as seen from Figure 13(c).
(a) Analytic form $\mathscr{K}$
(b) EinFields reconstructed $\hat{\mathscr{K}}$
@@ F.6.5 Kerr Kretschmann invariant
Figure 13: The Kretschmann scalar $\mathscr{K}$ of the Kerr metric computed in Cartesian Kerr-Schild form (Eq. (81)) in the $x$-$y$ plane for $z=0.3$.
F.7 Training on varied coordinate systems